OpenStudy (anonymous):

Someone help me solve these equations for x and y please? I need to show work too. Back breaking load of homework. -7x-3y=-25 -6x-4y=-20 Medal to first correct answer :)

OpenStudy (anonymous):

-7x-3y=-25 +3y +3y -7x=-25+3y divide both sides by -7 x=(25-3y)/7 -7x-3y=-25 +7x +7x -3y=7x-25 divide both sides by -3 y=(-7x+25)/3 -6x-4y=-20 +4y +4y -6x=4y-20 divide both sides by -6 x=(20-4y)/-6 -6x-4y=-20 +6x +6x -4y=6x-20 divide both sides by -4 y=5-(6/4)x

OpenStudy (anonymous):

so now i solve for y=5-(6/4)x?

OpenStudy (anonymous):

no that is the answer

OpenStudy (anonymous):

oh i get it, nvm thanks :D

OpenStudy (anonymous):

np